OpenStudy (anonymous):

Simplify. sqrt 48x^9

OpenStudy (anonymous):

\[\sqrt{48x^9}\]

OpenStudy (anonymous):

48 = 12* 4 = 4*4*3 \[\sqrt{4*4*3* x *x^4* x^4}\] now pull out those term which are 2 in number

OpenStudy (midhun.madhu1987):